OpenStudy (anonymous):

Find the inverse of the function. f(x) = 6x3 + 3

OpenStudy (jdoe0001):

so you find the inverse of a function by doing a quick "switcharoo" on the letters/variables so if the function is \(f(x) = 6x^3 + 3 \implies y = 6x^3 + 3\) the inverse will be by switching it about \(x = 6y^3 + 3 \) and solve for "y" there whatever you'd get, is the inverse function :)

OpenStudy (anonymous):

To find the inverse of the function f we write : \[f(y)=x\] and then we try to find the y ! We have : \[f(y)=x\iff 6y^3+3=x\\\iff 6y^3=x-3\\ \iff y^3=\frac{x-3}{6} \iff y=\sqrt[3]{\frac{x-3}6} \] So the inverse function of f is the function : \[f^{-1}(x)=\sqrt[3]{\frac{x-3}6}\]

OpenStudy (anonymous):

thank you